# Normalizing Flow-Assisted Nested Sampling on Type-II Seesaw Model

### v1

@@ Abstract
We propose a novel technique for sampling particle physics model parameter space. The main sampling method applied is Nested Sampling (NS), which is boosted by the application of multiple Machine Learning (ML) networks, e.g., Self-Normalizing Network (SNN) and Normalizing Flow (specifically RealNVP). We apply this on Type-II Seesaw model to test the efficacy of the algorithm. We present the results of our detailed Bayesian exploration of the model parameter space subjected to theoretical constraints and experimental data corresponding to the 125 GeV Higgs boson, $\rho$-parameter, and the oblique parameters.
Resources
@@
Associated data, figures, and trained ML models: GitHub [Uncaptioned image]
preprint: APS/123-QED
I Introduction
@@
Even as the large hadron collider (LHC) approaches its high luminosity era, a smoking gun signal for any beyond the standard model (BSM) physics scenario continues to elude us. However, we have a wealth of experimental data not only from the LHC but from other particle physics experiments as well, such as dark matter, neutrino oscillation, lepton flavor violating (LFV) and lepton number violating (LNV) decay search experiments, etc. As a result, phenomenological BSM searches have shifted more towards constraining new physics (NP) domains using the available data rather than predicting rates of new experimental observables by choosing random benchmark points from different regions of parameter space (PS). Even if one takes the second approach, one must have a clear idea about the available parameter space, consistent with existing data. This becomes increasingly difficult with the increasing dimensionality of NP parameter space. A large number of input parameters coupled with the growing precision of data makes the parameter space of most BSM scenarios intractable.
@@
The most common statistically viable way of handling this problem is a likelihood-based estimation (the likelihood is computed with the help of experimental data on a set of relevant observables), where one either finds the best-fit point in the PS with the associated confidence intervals by optimizing the negative log-likelihood or calculates the posterior distribution of the parameters. The most commonly used method to find the (unnormalized) posterior is Markov Chain Monte Carlo (MCMC), with all of its variants [1, 2, 3], which, while not suitable to find the exact evidence in favor of the data, supplies a sample from the posterior. This enables one to plot one- or two-dimensional marginal credible interval contours from the posterior. These tell us which part of the parameter space one should concentrate upon for any phenomenological study. However, a well-known issue with MCMC is its tendency to get stuck in one mode of the PS, provided it finds one [4]. This becomes problematic for multi-modal spaces typical for many particle physics models [5, 6, 7]. As a Markov chain, MCMC is, in theory, capable of finding the other modes, but takes an a priori unknown amount of time in the process. A single chain is not parallelizable and methods involving multiple chains [8, 9] increase the computational cost. Convergence criteria for these chains are neither foolproof nor beyond debate [10]. The scenario worsens in the presence of typical theoretical constraints which prevent high-energy phenomenological (HEP) spectrum generators to generate consistent spectrum in a large chuck of the parameter space. All of these, in aggregate, make MCMC computationally expensive for complicated BSM parameter spaces. Application of machine learning (ML) algorithms to improve upon the existing techniques in this regard has gained relevance in recent times [11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21]. For a more comprehensive account of these works one may refer to [22].
@@
Newer, more efficient statistical techniques such as Nested Sampling (NS) [23, 24] (auto-tuned convergence, access to multiple modes, parallelizability) are needed to circumvent the issues MCMC has. NS is a numerical integration strategy that can be used to obtain Bayesian evidence and, in the process, one gets a weighted sample from the posterior. A well-known HEP package GAMBIT [25, 26] uses MultiNest [27, 28] to perform NS on BSM parameter space. However, as the algorithm moves toward high-likelihood regions, the progress becomes slower (and convergence takes longer) because of the rarity of higher likelihood points. We propose an ML solution to this. A special type of artificial neural network (ANN) is trained as a classifier to circumnavigate the external constraints, and a trained Normalizing Flow (NF)-type network is used to reduce the convergence time. The latter network is trained iteratively during the running of NS, such that, by the time NS reaches high-likelihood regions, it is trained well enough to predict a much larger number of higher likelihood points than a naive sampling of the prior.
@@
We have adopted a well-motivated BSM scenario, namely, the Type-II seesaw model to showcase the efficacy of this methodology. Type-II seesaw mechanism [29, 30, 31, 32, 33, 34] provides us with one of the simplest frameworks to generate light neutrino masses and mixing angles [35, 36, 37, 38, 38, 39, 40] without introducing any right-handed neutrinos into the theory. Instead, this model introduces a complex scalar, a triplet under the ${\rm SU(2)_{L}}$ gauge group. A Majorana mass term is generated for the light neutrinos after electroweak symmetry breaking through an LNV Yukawa coupling and the triplet vacuum expectation value (VEV). The model has very rich phenomenological implications [41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81] owing to the exotic scalar sector and neutrino mass generation. The scalar sector consists of one doubly charged scalar, one singly charged scalar, and three neutral scalars: two CP-even and one CP-odd. In addition to that, lepton number violating interactions lead to predictions of lepton number and lepton flavor violating decays. All these factors combine to make Type-II seesaw a highly motivated and phenomenologically interesting scenario. The triplet VEV ($v_{t}$) is instrumental in this context. A sufficiently small $v_{t}$ ($\lesssim 10^{-9}$) allows the lepton number violating Yukawa couplings to be large enough in the context of light neutrino mass and mixing, but that also in turn leads to stringent constraints arising from non-observation of lepton number and/or lepton flavor violation. Such small $v_{t}$, however, does not affect electroweak precision data such as $W$ and $Z$ boson masses, $\rho$ parameter, etc. In this work, we explore the large $v_{t}$ scenario and ascertain how the parameter space shapes up when confronted with the electroweak precision data [82, 83, 84, 85, 86, 87, 88] and the 125 GeV Higgs data [89, 90, 91, 92, 93, 94]. We also estimate the allowed numerical ranges of the LNV Yukawa parameters subjected to the neutrino oscillation data [40]. Although this scenario has been explored before, a posterior distribution with numerical estimates of central tendency and dispersion for all the new parameters introduced in this model is presented here for the first time.
@@
The article is organized as follows. Section II briefly describes the Type-II Seesaw model. Section III starts with a discussion about the observable set and parameter ranges used in the analysis. Then we give a detailed account of the algorithm: the ML networks and their performance in particular. In section IV, after presenting the parameter posterior, we proceed to discuss the properties of the scalar particles in the model within the favored parameter space and their collider constraints. We conclude in section V.
II Theory
@@
Here $\sigma^{2}$ indicates the second Pauli matrix. After electroweak symmetry breaking $\Phi$ and $\Delta$ acquire VEVs $\frac{v_{d}}{\sqrt{2}}$ and $\frac{v_{t}}{\sqrt{2}}$ respectively, with the total electroweak VEV given by $v=\sqrt{v_{d}^{2}+2v_{t}^{2}}$. The parameters $m^{2}_{\Phi}$ and $m^{2}_{\Delta}$ are not independent. They can be determined in terms of other free parameters using the tadpole equations:
@@
In the scalar sector, the neutral CP-even states can have considerably large mixing depending on the choices of new physics parameters. The mass-squared matrix at the tree level can be written as
@@
However, large mixing between the CP-even states is disfavored from the 125 GeV Higgs data. The mixing angle ($\alpha$) is given as
@@
The five quartic couplings introduced in the scalar potentials, namely, $\lambda$, $\lambda_{1}$, $\lambda_{2}$, $\lambda_{3}$ and $\lambda_{4}$ are strictly not independent of each other. Their choices are restricted theoretically from the requirements of vacuum stability and perturbative unitarity. These constraints have been studied in detail in the context of the Type-II seesaw model [95, 58]. For the constraint equations applicable to our choice of quartic coupling assignment, please refer to [81].
@@
III Methodology
III.1 Observables and Parameters
The model is implemented in the Mathematica package SARAH [96, 97, 98] which performs analytical calculations about mass matrices and vertices up to one loop. Full two-loop corrections are implemented for the scalar sector [99, 100]. SARAH also generates source codes for SPheno [101, 102] based on these analytical calculations. SPheno performs the numerical computation to generate a full particle spectrum including masses, mixing, and decay information of all particles. SPheno also provides the platform to check low-energy flavor constraints through its interface with FlavorKit [103]. We have used the spectrum file generated by SPheno to check constraints from the SM-like Higgs data through HiggsBounds (HB) [104, 105, 106, 107, 108, 109]. HB determines whether a parameter point is excluded or not at $95\%$ confidence level (CL) based on the scalar search data obtained from LEP, Tevatron and the LHC. For likelihood calculation of the scalar sector at each parameter point we have used HiggsSignals (HS) [106, 110, 111, 112, 113]. HS provides a peak-centered $\chi^{2}$ measure of compatibility between the measured (SM-like) Higgs signal rates and masses obtained from the LHC and Tevatron and the model predictions for a fixed (hypothetical) Higgs mass. The combination of these two packages was used in our analysis as a C++ implementation HiggsTools [114].
We have considered 163 observables in total: the full set of 159 Higgs data used by HiggsSignals, the three oblique parameters and the $\rho$-parameter. The Higgs sector observables are calculated by HS using the spectrum generated by SPheno. The rest of the observables are computed up to one-loop level by SPheno. The oblique parameter fits are performed by the GFitter collaboration [86] and we use the data and correlation from there. The $\rho$ parameter data is obtained from [87, 88].
@@ III.1 Observables and Parameters
III.2 Framework
III.2.1 Set-up
We aim to perform an NS over the parameter space, to obtain a Bayesian posterior distribution dictated by the data used. Details about NS can be found in ref.s [23, 24, 115, 22]. In short, the algorithm starts with $n_{Live}$live’ points sampled from a prior over the parameter space, generates more independent and identically distributed (iid) points in a specified region until it finds one point with likelihood ($\mathcal{L}$) greater than the lowest likelihood among the live points ($\mathcal{L^{*}}$), replaces that lowest-likelihood point with the new one, saves the replaced point as a weighted sample, shrinks the target region in a specified way and then repeats. This way, evaluating a multi-dimensional integral essentially changes to a single-variable one, by peeling off equal-probability-contours from the parameter space. Mainly an algorithm to calculate the evidence/marginal likelihood of the posterior, NS generates a weighted sample drawn from the posterior, as a byproduct. In addition to being self-tuning, mostly parallelizable, and able to sample from the whole prior space simultaneously—thus handling multi-modal problems much better than its MCMC counterparts—NS is also much faster than other equivalent MCMC methods while spending less computational resources. Due to the way NS works, it gets slower, the closer it gets to the high-likelihood region, especially if the shrinkage of the sampling region is comparatively less.
The main bottleneck in this kind of work is calculating the complicated likelihood function repetitively for different parameter points. Another additional source of slowdown is that not all points are allowed in the parameter space. Some of them are discarded due to directly applicable (thus quickly calculable) theoretical constraints; others, though allowed by the constraints mentioned, generate no spectrum (by some HEP spectrum generator like SPheno, mainly due to unstable vacuum and tachyonic masses in the spectrum). ML, especially generative algorithms, can provide speed-ups here. In addition to the application of a classifier to predict the valid points, the most obvious solution to learn the data-consistent parameter space is to learn the structure of the optimization criteria (negative log-likelihood or $\chi^{2}$ in our context) shared by similar points in the parameter space, known as ‘Amortized Optimization’ elsewhere in the literature [116]. The prediction uncertainty notwithstanding, this approach has numerous problems:

### v2

@@ Abstract
We propose a novel technique for sampling particle physics model parameter space. The main sampling method applied is Nested Sampling (NS), which is boosted by the application of multiple Machine Learning (ML) networks, e.g., Self-Normalizing Network (SNN) and Normalizing Flow (specifically RealNVP). We apply this to the Type-II Seesaw model to test the algorithm’s efficacy. We present the results of our detailed Bayesian exploration of the model parameter space subjected to theoretical constraints and experimental data corresponding to the 125 GeV Higgs boson, $\rho$-parameter, and the oblique parameters.
@@
All associated data, figures, and trained ML models can be found here: GitHub [Uncaptioned image]
pacs:
PACS-keydiscribing text of that key and PACS-keydiscribing text of that key
1 Introduction
@@
Even as the large hadron collider (LHC) approaches its high luminosity era, a smoking gun signal for any beyond the standard model (BSM) physics scenario continues to elude us. However, we have a wealth of experimental data not only from the LHC but from other particle physics experiments as well, such as dark matter, neutrino oscillation, lepton flavour violating (LFV) and lepton number violating (LNV) decay search experiments, etc. As a result, phenomenological BSM searches have shifted more towards constraining new physics (NP) domains using the available data rather than predicting rates of new experimental observables by choosing random benchmark points from different regions of parameter space (PS). Even if one takes the second approach, one must have a clear idea about the available parameter space, consistent with existing data. This becomes increasingly difficult with the increasing dimensionality of NP parameter space. A large number of input parameters coupled with the growing precision of data makes the parameter space of most BSM scenarios intractable.
@@
The most common statistically viable way of handling this problem is a likelihood-based estimation (the likelihood is computed with the help of experimental data on a set of relevant observables), where one either finds the best-fit point in the PS with the associated confidence intervals by optimizing the negative log-likelihood or calculates the posterior distribution of the parameters. The most commonly used method to find the (unnormalized) posterior is Markov Chain Monte Carlo (MCMC), with all of its variants Hastings:1970 ; Speagle:2019ffr ; brooks2011eds , which, while not suitable to find the exact evidence in favour of the data, supplies a sample from the posterior. This enables one to plot one- or two-dimensional marginal credible interval contours from the posterior. These tell us which part of the parameter space one should concentrate upon for any phenomenological study. However, a well-known issue with MCMC is its tendency to get stuck in one mode of the PS, provided it finds one tripuraneni2017magnetic . This becomes problematic for multi-modal spaces typical for many particle physics models Albert:2024zsh ; GAMBIT:2017zdo ; Kvellestad:2017rwl . As a Markov chain, MCMC is, in theory, capable of finding the other modes, but takes an a priori unknown amount of time in the process. A single chain is not parallelizable and methods involving multiple chains foreman2013emcee ; DeBlas:2019ehy increase the computational cost. Convergence criteria for these chains are neither foolproof nor beyond debate Cowles01081999 . The scenario worsens in the presence of typical theoretical constraints which prevent high-energy phenomenological (HEP) spectrum generators to generate consistent spectrum in a large chuck of the parameter space. All of these, in aggregate, make MCMC computationally expensive for complicated BSM parameter spaces. Application of machine learning (ML) algorithms to improve upon the existing techniques in this regard has gained relevance in recent times Caron:2019xkx ; Hollingsworth:2021sii ; DBLP:journals/corr/abs-2107-03342 ; Hammad:2022wpq ; Goodsell:2022beo ; Diaz:2024sxg ; Binjonaid:2024jpm ; Hammad:2024tzz ; Hermans:2019ioj ; Brehmer:2020cvb ; Morrison:2022vqe ; Hunt-Smith:2023ccp ; Albergo:2021vyo . For a more comprehensive account of these works one may refer to Baruah:2024gwy .
@@
Newer, more efficient statistical techniques such as Nested Sampling (NS) Skilling:2004 ; Skilling:2006 (auto-tuned convergence, access to multiple modes, parallelizability) are needed to circumvent the issues MCMC has. NS is a numerical integration strategy that can be used to obtain Bayesian evidence and, in the process, one gets a weighted sample from the posterior. A well-known HEP package GAMBIT GAMBIT:2017yxo ; Bloor:2021gtp uses MultiNest Feroz:2008xx ; Feroz:2009 to perform NS on BSM parameter space. However, as the algorithm moves toward high-likelihood regions, the progress becomes slower (and convergence takes longer) because of the rarity of higher likelihood points. We propose an ML solution to this. A special type of artificial neural network (ANN) is trained as a classifier to circumnavigate the external constraints, and a trained Normalizing Flow (NF)-type network is used to reduce the convergence time. The latter network is trained iteratively during the running of NS, such that, by the time NS reaches high-likelihood regions, it is trained well enough to predict a much larger number of higher likelihood points than a naive sampling of the prior.
@@
We have adopted a well-motivated BSM scenario, namely, the Type-II seesaw model to showcase the efficacy of this methodology. Type-II seesaw mechanism Konetschny:1977bn ; Magg:1980ut ; Schechter:1980gr ; Lazarides:1980nt ; Mohapatra:1980yp ; Cheng:1980qt provides us with one of the simplest frameworks to generate light neutrino masses and mixing angles RENO:2018dro ; DayaBay:2018yms ; Super-Kamiokande:2019gzr ; T2K:2018rhz ; T2K:2018rhz ; NOvA:2019cyt ; deSalas:2020pgw without introducing any right-handed neutrinos into the theory. Instead, this model introduces a complex scalar, a triplet under the ${\rm SU(2)_{L}}$ gauge group. A Majorana mass term is generated for the light neutrinos after electroweak symmetry breaking through an LNV Yukawa coupling and the triplet vacuum expectation value (VEV). The model has very rich phenomenological implications Huitu:1996su ; Chakrabarti:1998qy ; Muhlleitner:2003me ; Akeroyd:2005gt ; Perez:2008ha0 ; delAguila:2008cj ; Akeroyd:2009hb ; Melfo:2011nx ; Arhrib:2011vc ; Akeroyd:2012nd ; Chun:2012zu ; Chun:2012jw ; Englert:2013wga ; Kanemura:2013vxa ; Chun:2013vma ; Chun:2003ej ; Dey:2008jm ; Arhrib:2011uy ; Dev:2013ff ; Arhrib:2014nya ; Das:2016bir ; Anisha:2021fzf ; Du:2018eaw ; Li:2018jns ; Kanemura:2014goa ; Han:2015hba ; Han:2015sca ; Mitra:2016wpr ; Babu:2016rcr ; BhupalDev:2018tox ; Antusch:2018svb ; Ferreira:2019qpf ; Banerjee:2024jwn ; Ashanujjaman:2023tlj ; Giarnetti:2023dcr ; Chiang:2021lsx ; Ashanujjaman:2022ofg ; Ashanujjaman:2022tdn ; Ashanujjaman:2021txz ; Ducu:2024xxf ; Primulando:2019evb ; Das:2023tna ; Das:2024yvt ; Das:2024kyk owing to the exotic scalar sector and neutrino mass generation. The scalar sector consists of one doubly charged scalar, one singly charged scalar, and three neutral scalars: two CP-even and one CP-odd. In addition to that, lepton number-violating interactions lead to predictions of lepton number and lepton flavour-violating decays. All these factors combine to make Type-II seesaw a highly motivated and phenomenologically interesting scenario. The triplet VEV ($v_{T}$) is instrumental in this context. A sufficiently small $v_{T}$ ($\lesssim 10^{-9}$) allows the lepton number violating Yukawa couplings to be large enough in the context of light neutrino mass and mixing, but that also in turn leads to stringent constraints arising from non-observation of lepton number and/or lepton flavour violation. Such small $v_{T}$, however, does not affect electroweak precision data such as $W$ and $Z$ boson masses, $\rho$ parameter, etc. In this work, we explore the large $v_{T}$ scenario and ascertain how the parameter space shapes up when confronted with the electroweak precision data Amaldi:1987fu ; Costa:1987qp ; Langacker:1991an ; Erler:1994fz ; Haller:2018nnx ; ParticleDataGroup:2022pth ; ParticleDataGroup:2024cfk and the 125 GeV Higgs data CMS:2012qbp ; ATLAS:2012yve ; ATLAS:2018tdk ; CMS:2020xrn ; ATLAS:2021vrm ; cms_web1 . We also estimate the allowed numerical ranges of the LNV Yukawa parameters subjected to the neutrino oscillation data deSalas:2020pgw . Although this scenario has been explored before, a posterior distribution with numerical estimates of central tendency and dispersion for all the new parameters introduced in this model is presented here for the first time.
@@
The article is organized as follows. Section 2 briefly describes the Type-II Seesaw model. Section 3 starts with a discussion about the observable set and parameter ranges used in the analysis. Then we give a detailed account of the algorithm: the ML networks and their performance in particular. In section 4, after presenting the parameter posterior, we proceed to discuss the properties of the scalar particles in the model within the favoured parameter space and their collider constraints. We conclude in section 5.
2 Theory
@@
Here $\sigma^{2}$ indicates the second Pauli matrix. $\mu_{1}$ is a trilinear coupling, whereas, $\lambda$ and $\lambda_{i}~{}(i=1,2,3,4)$ represent the quartic couplings. After electroweak symmetry breaking $\Phi$ and $\Delta$ acquire VEVs $\frac{v_{d}}{\sqrt{2}}$ and $\frac{v_{T}}{\sqrt{2}}$ respectively, with the total electroweak VEV given by $v=\sqrt{v_{d}^{2}+2v_{T}^{2}}$. The parameters $m^{2}_{\Phi}$ and $m^{2}_{\Delta}$ are not independent. They can be determined in terms of other free parameters using the tadpole equations:
@@
In the scalar sector, the neutral CP-even states can have considerably large mixing depending on the choices of new physics parameters. The mass-squared matrix at the tree level can be written as {strip}
@@
   However, a large mixing between the CP-even states is disfavoured from the 125 GeV Higgs data. The mixing angle ($\alpha$) is given as
@@
The five quartic couplings introduced in the scalar potentials, namely, $\lambda$, $\lambda_{1}$, $\lambda_{2}$, $\lambda_{3}$ and $\lambda_{4}$ are strictly not independent of each other. Their choices are restricted theoretically from the requirements of vacuum stability and perturbative unitarity. These constraints have been studied in detail in the context of the Type-II seesaw model Bonilla:2015eha ; Arhrib:2011uy . For the constraint equations applicable to our choice of quartic coupling assignment, please refer to Primulando:2019evb .
@@
3 Methodology
3.1 Observables and Parameters
The model is implemented in the Mathematica package SARAH [^1] Version: 4.14.4 Staub:2008uz ; Staub:2010jh ; Staub:2015kfa which performs analytical calculations about mass matrices and vertices up to one loop. Full two-loop corrections are implemented for the scalar sector Goodsell:2014bna ; Goodsell:2015ira . SARAH also generates source codes for SPheno Porod:2003um ; Porod:2011nf based on these analytical calculations. SPheno performs the numerical computation to generate a full particle spectrum including masses, mixing, and decay information of all particles. SPheno also provides the platform to check low-energy flavour constraints through its interface with flavourKit Porod:2014xia . We have used the spectrum file generated by SPheno [^2] Version: 4.0.4 to check constraints from the SM-like Higgs data through HiggsBounds (HB) Bechtle:2008jh ; Bechtle:2011sb ; Bechtle:2012lvg ; Bechtle:2013wla ; Bechtle:2015pma ; Bechtle:2020pkv . HB determines whether a parameter point is excluded or not at $95\%$ confidence level (CL) based on the scalar search data obtained from LEP, Tevatron and the LHC. For likelihood calculation of the scalar sector at each parameter point we have used HiggsSignals (HS) Bechtle:2012lvg ; Bechtle:2013xfa ; Stal:2013hwa ; Bechtle:2014ewa ; Bechtle:2020uwn . HS provides a peak-centered $\chi^{2}$ measure of compatibility between the measured (SM-like) Higgs signal rates and masses obtained from the LHC and Tevatron and the model predictions for a fixed (hypothetical) Higgs mass. The combination of these two packages was used in our analysis as a C++ implementation HiggsTools [^3] Combination of HiggsBounds-5 and HiggsSignals-2 Bahl:2022igd .
We have considered 163 observables in total: the full set of 159 Higgs data used by HiggsSignals, the three oblique parameters and the $\rho$-parameter. The Higgs sector observables are calculated by HS using the spectrum generated by SPheno. The rest of the observables are computed up to one-loop level by SPheno. The oblique parameter fits are performed by the GFitter collaboration Haller:2018nnx and we use the data and correlation from there. The $\rho$ parameter data is obtained from ParticleDataGroup:2022pth ; ParticleDataGroup:2024cfk .
@@ 3.1 Observables and Parameters
3.2 Framework
3.2.1 Set-up
We aim to perform an NS over the parameter space, to obtain a Bayesian posterior distribution dictated by the data used. Details about NS can be found in ref.s Skilling:2004 ; Skilling:2006 ; Ashton:2022 ; Baruah:2024gwy . In short, the algorithm starts with $n_{Live}$live’ points sampled from a prior over the parameter space, generates more independent and identically distributed (iid) points in a specified region until it finds one point with likelihood ($\mathcal{L}$) greater than the lowest likelihood among the live points ($\mathcal{L^{*}}$), replaces that lowest-likelihood point with the new one, saves the replaced point as a weighted sample, shrinks the target region in a specified way and then repeats. This way, evaluating a multi-dimensional integral essentially changes to a single-variable one, by peeling off equal-probability-contours from the parameter space. Mainly an algorithm to calculate the evidence/marginal likelihood of the posterior, NS generates a weighted sample drawn from the posterior, as a by-product. In addition to being self-tuning, mostly parallelizable, and able to sample from the whole prior space simultaneously—thus handling multi-modal problems much better than its MCMC counterparts—NS is also much faster than other equivalent MCMC methods while spending less computational resources. Due to the way NS works, it gets slower, the closer it gets to the high-likelihood region, especially if the shrinkage of the sampling region is comparatively less.
The main bottleneck in this kind of work is calculating the complicated likelihood function repetitively for different parameter points. Another additional source of slowdown is that not all points are allowed in the parameter space. Some of them are discarded due to directly applicable (thus quickly calculable) theoretical constraints; others, though allowed by the constraints mentioned, generate no spectrum (by some HEP spectrum generator like SPheno, mainly due to unstable vacuum and tachyonic masses in the spectrum). ML, especially generative algorithms, can provide speed-ups here. In addition to the application of a classifier to predict the valid points, the most obvious solution to learn the data-consistent parameter space is to learn the structure of the optimization criteria (negative log-likelihood or $\chi^{2}$ in our context) shared by similar points in the parameter space, known as ‘Amortized Optimization’ elsewhere in the literature amos2023tutorial . The prediction uncertainty notwithstanding, this approach has numerous problems: